Silver Company makes a product that is very popular as a Mother’s Day gift. Thus, peak sales occur in May of each year, as shown in the company’s sales budget for the second quarter given below:

April May June Total
Budgeted sales (all on account) $320,000 $520,000 $180,000 $1,020,000

From past experience, the company has learned that 30% of a month’s sales are collected in the month of sale, another 65% are collected in the month following sale, and the remaining 5% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $250,000, and March sales totaled $280,000.

Required:

1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

2. What is the accounts receivable balance on June 30th?
Answer #1
1
April May June Total
February sales 12500 12500
March sales 182000 14000 196000
April sales 96000 208000 16000 320000
May sales 156000 338000 494000
June Sales 54000 54000
Total cash collections 290500 378000 408000 1076500
2
May sales 26000 =520000*5%
June Sales 126000 =180000*70%
Accounts receivable balance on June 30 152000
Workings:
April May June
February sales =250000*5%
March sales =280000*65% =280000*5%
April sales =320000*30% =320000*65% =320000*5%
May sales =520000*30% =520000*65%
June Sales =180000*30%